# Multi-View Spectrogram Transformer for Respiratory Sound Classification

### v2

@@ Abstract
Deep neural networks have been applied to audio spectrograms for respiratory sound classification. Existing models often treat the spectrogram as a synthetic image while overlooking its physical characteristics. In this paper, a Multi-View Spectrogram Transformer (MVST) is proposed to embed different views of time-frequency characteristics into the vision transformer. Specifically, the proposed MVST splits the mel-spectrogram into different-sized patches, representing the multi-view acoustic elements of a respiratory sound. These patches and positional embeddings are then fed into transformer encoders to extract the attentional information among patches through a self-attention mechanism. Finally, a gated fusion scheme is designed to automatically weigh the multi-view features to highlight the best one in a specific scenario. Experimental results on the ICBHI dataset demonstrate that the proposed MVST significantly outperforms state-of-the-art methods for classifying respiratory sounds.
@@ 1 Introduction
Fig. 1: Overview of proposed Multi-View Spectrogram Transformer (MVST). The mel spectrogram of an input audio is split into different-sized patches, to analyze the multi-view spectral characteristics. The multi-view spectrogram patches are then processed by multi-scale transformer encoders to capture the attentional information among patches. A gated fusion mechanism is then designed to highlight the most suitable spectral view for respiratory sound classification.
@@
The proposed MSVT is motivated by two factors. 1) The mel spectrogram provides fine resolution for lower frequencies but coarser resolution for higher frequencies [23], while the square patches in audio spectrograms [21, 22] often results in limited spectral information of high-frequency patches. 2) The same type of respiratory sounds may contain significant frequency changes [12], which leads to misalignment of patches in frequency if the patch is too small, or losing the critical frequency information if the patch is too large. To address these two challenges, the proposed MVST embeds different views of spectral characteristics into ViTs using a novel patching scheme that treats the time and frequency axes differently. Specifically, the input mel spectrogram is split into patches of shape varying from $16\times 16$ as in the ViT [16] to $256\times 1$ by increasing the bandwidth while reducing the time interval. This partially solves the issue of frequency misalignment. More importantly, patches of a specific size capture a unique view of the mel spectrogram, and collectively these multi-view spectrograms capture different spectral views of a given audio signal. All the patches are then embedded with positional information and fed into the transformer encoders to perform multi-head self-attention, which deeply exploits the spectral characteristics of the audio signal. A gated fusion scheme [24] is then designed to automatically highlight the most suitable view of spectral features for final classification.
@@ 2.1 Overview of Proposed MVST
The proposed MVST for respiratory sound classification is shown in Fig. 1. The raw audio signal is first converted into a mel spectrogram using STFT. To capture the multi-view spectral characteristics embedded in different frequency bands at different time intervals, the spectrogram is split into different-sized patches as described in Sec. 2.2. These patches together with the position embeddings are then fed into the transformer encoders, in which the multi-head self-attention (MSA) [16] is utilized to exploit the audio characteristics residing in patches and the long-range interactions between patches as described in Sec. 2.3. As a result, patches with bare spectral information will be less weighed while patches with rich spectral information will be weighed with a greater attentive importance. The recurrent spectral patterns can be more effectively captured across patches. Lastly, the spectral features of different views are fused through the gated fusion scheme for final classification as described in Sec. 2.4.
@@
A linear transformation is then applied on the spectrogram to convert it into mel scale to derive the mel spectrogram.
@@ 2.2 Multi-View Patch Splitting
The spectrogram $\bm{S}$ is first resized to $N\times N$. The multi-view patch splitting then divides it into patches of size $\mu\times\nu$ to obtain the set of patches $\bm{\mathsf{P}}\in\mathbb{R}^{\frac{N}{\mu}\times\frac{N}{\nu}\times\mu\nu}$, which represent $n_{T}=\frac{N}{\mu}\frac{N}{\nu}$ token vectors of dimension $\mu\nu$. In the standard ViT [16], a common choice is a square with $\mu=\nu=16$. The proposed multi-view patch splitting generates patches of different shapes from the spectrogram by considering the time and frequency differently, e.g., for a patch of size $2^{\ell}\times\frac{N}{2^{\ell}}$, $\bm{\mathsf{P}}^{\ell}\in\mathbb{R}^{\frac{N}{2^{\ell}}\times 2^{\ell}\times N}$. Each group of patches represent a unique view of audio spectral characteristics, e.g., square patches represent the spectral information in a specific frequency band at a specific time interval while slim and long patches of size $N\times 1$ represent the entire frequency response at a specific time instance. In practice, five groups of patches $\{{\bm{\mathsf{P}}}^{\ell},\ell=0\dots 4\}$ are generated from the spectrogram of size $256\times 256$, whose size ranges from $256\times 1$, $128\times 2$, $64\times 4$, $32\times 8$ to $16\times 16$, as shown in Fig. 1.
Fig. 2: Two example spectrograms sliced from the same crackling breathing cycle at different time intervals.
@@ 3.1 Experimental Settings
The proposed MVST is evaluated on the largest publicly available respiratory sound dataset, the ICBHI dataset [1], which contains 6,898 respiratory cycles, including 3,642 normal breathing, 1,864 crackling breathing, 886 wheezing breathing and 506 breathing with both crackling & wheezing. The sampling rate is 4kHz, 10kHz or 44.1kHz, with a recording duration between 10 and 90 seconds. Each recording includes an average of seven breathing cycles. Following the official experimental settings in [1, 2], 60% of the recordings are randomly selected for training and the remaining 40% for testing. Standard evaluation metrics [1] such as specificity (SP), sensitivity (SE) and average score (AS) are used. For the proposed MVST, the cross-entropy loss function is used with an initial learning rate of $1\times 10^{-3}$. The AdamW optimizer is used with a fixed weight decay of $1\times 10^{-5}$. To ensure the convergence and the efficiency of the model, the maximum number of epochs is set to $100$ and the batch size is $64$.
@@ 3.2 Ablation Study
An ablation study is carried out to evaluate the effectiveness of the proposed MVST. For the five compared single-view methods, the spectrograms are divided into patches of the specific size and fed into the ViT for classification. The results are summarized in Table 1. It can be observed that: 1) Many other views like $32\times 8$, $64\times 4$ and $128\times 2$ perform better than the $16\times 16$ baseline method, which shows the limitation of the baseline method. 2) The $128\times 2$ patching scheme produces the highest AS of 63.91% among single-view methods, which demonstrates the effectiveness of generating slim and long patches by reducing the time interval while increasing the bandwidth, as this design partially alleviates the problem of frequency misalignment in respiratory sounds. 3) The proposed MVST produces significantly better SP, SE and AS than all single-view methods, which is attributed to the multi-view spectrograms in effectively capturing the audio characteristics, and the gated fusion scheme in combining the multi-view spectrogram features for robust respiratory classification.
@@
$32\times 8$ 78.55 47.50 63.02
$64\times 4$ 79.70 46.10 62.90
$128\times 2$ 78.12 49.71 63.91
$256\times 1$ 80.92 43.75 62.33
@@ 3.3 Comparison to State-of-the-art Methods
Many deep neural networks have been utilized as the backbone for respiratory sound classification, e.g., bi-ResNet in LungBRN [4], ResNet-Att in LungAttn [5], ResNeSt [7], ResNet-34 in RespireNet [6], bi-ResNet-Att in ARSC-Net [3], CNN6 [10], Co-tuned ResNet-50 [2] and audio spectrogram transformer with patch-mix and contrastive learning (AST + patch-mix CL[22]. The comparison with these state-of-the-art methods are summarized in Table 2.111The results of all the compared methods are obtained from here. It can be observed that: 1) Many models utilize traditional CNN models for respiratory classification [2, 4, 5, 6, 3, 7, 10], while the most recent work [22] utilizes audio spectrogram transformer [21] and achieves the previous best results on the ICBHI dataset. This partially motivates us to utilize the ViT as the backbone network. 2) The proposed MVST significantly outperforms all the compared methods in terms of all three criteria SP, SE and AS. Specifically, it outperforms the second-best approach AST + patch-mix CL [22] by 0.33%, 8.09% and 4.18% in terms of SP, SE and AS respectively. Compared to the second best results by Co-tuned ResNet-50 [2], the proposed MVST achieves a performance gain as high as 8.26% in terms of the key metric AS.

### v3

@@ Abstract
Deep neural networks have been applied to audio spectrograms for respiratory sound classification. Existing models often treat the spectrogram as a synthetic image while overlooking its physical characteristics. In this paper, a Multi-View Spectrogram Transformer (MVST) is proposed to embed different views of time-frequency characteristics into the vision transformer. Specifically, the proposed MVST splits the mel-spectrogram into different-sized patches, representing the multi-view acoustic elements of a respiratory sound. The patches and positional embeddings are fed into transformer encoders to extract the attentional information among patches through a self-attention mechanism. Finally, a gated fusion scheme is designed to automatically weigh the multi-view features to highlight the best one in a specific scenario. Experimental results on the ICBHI dataset demonstrate that the MVST significantly outperforms state-of-the-art methods for classifying respiratory sounds. The code is available at: https://github.com/wentaoheunnc/MVST.
@@ 1 Introduction
Fig. 1: Overview of proposed Multi-View Spectrogram Transformer (MVST). The mel-spectrogram of an input audio is split into different-sized patches, to analyze the multi-view spectral characteristics. The multi-view spectrogram patches are then processed by multi-scale transformer encoders to capture the attentional information among patches. A gated fusion mechanism is then designed to highlight the most suitable spectral view for respiratory sound classification.
@@
The proposed MSVT is motivated by two factors. 1) The mel-spectrogram provides fine resolution for lower frequencies but coarser resolution for higher frequencies [23], while the square patches in audio spectrograms [21, 22] often results in limited spectral information of high-frequency patches. 2) The same type of respiratory sounds may contain significant frequency changes [12], which leads to misalignment of patches in frequency if the patch is too small, or losing the critical frequency information if the patch is too large. To address these two challenges, the proposed MVST embeds different views of spectral characteristics into ViTs using a novel patching scheme that treats the time and frequency axes differently. Specifically, the input mel-spectrogram is split into patches of shape varying from $16\times 16$ as in the ViT [16] to $256\times 1$ by increasing the bandwidth while reducing the time interval. This partially solves the issue of frequency misalignment. More importantly, patches of a specific size capture a unique view of the mel-spectrogram, and collectively these multi-view spectrograms capture different spectral views of a given audio signal. All the patches are then embedded with positional information and fed into the transformer encoders to perform multi-head self-attention, which deeply exploits the spectral characteristics of the audio signal. A gated fusion scheme [24] is then designed to automatically highlight the most suitable view of spectral features for final classification.
@@ 2.1 Overview of Proposed MVST
The proposed MVST for respiratory sound classification is shown in Fig. 1. The raw audio signal is first converted into a mel-spectrogram using STFT. To capture the multi-view spectral characteristics embedded in different frequency bands at different time intervals, the spectrogram is split into different-sized patches as described in Sec. 2.2. These patches together with the position embeddings are then fed into the transformer encoders, in which the multi-head self-attention (MSA) [16] is utilized to exploit the audio characteristics residing in patches and the long-range interactions between patches as described in Sec. 2.3. As a result, patches with bare spectral information will be less weighed while patches with rich spectral information will be weighed with a greater attentive importance. The recurrent spectral patterns can be more effectively captured across patches. Lastly, the spectral features of different views are fused through the gated fusion scheme for final classification as described in Sec. 2.4.
@@
A linear transformation is then applied on the spectrogram to convert it into mel scale to derive the mel-spectrogram.
@@ 2.2 Multi-View Patch Splitting
The spectrogram $\bm{S}$ is first resized to $H\times W$. The multi-view patch splitting then divides it into patches of size $\mu\times\nu$ to obtain the set of patches $\bm{\mathsf{P}}\in\mathbb{R}^{\frac{H}{\mu}\times\frac{W}{\nu}\times\mu\nu}$, which represent $n_{T}=\frac{H}{\mu}\frac{W}{\nu}$ token vectors of dimension $\mu\nu$. In the standard ViT [16], a common choice is a square with $\mu=\nu=16$. The proposed multi-view patch splitting generates non-overlapped patches of different shapes from the spectrogram by considering the time and frequency differently, e.g., for a patch of size $2^{\ell}\times\frac{H}{2^{\ell}}$, $\bm{\mathsf{P}}^{\ell}\in\mathbb{R}^{\frac{H}{2^{\ell}}\times 2^{\ell}\times H}$. Each group of patches represent a unique view of audio spectral characteristics, e.g., square patches represent the spectral information in a specific frequency band at a specific time interval while slim and long patches of size $H\times 1$ represent the entire frequency response at a specific time instance. In practice, five groups of patches $\{{\bm{\mathsf{P}}}^{\ell},\ell=0\dots 4\}$ are generated from the spectrogram of size $256\times 1024$, whose size ranges from $256\times 1$, $128\times 2$, $64\times 4$, $32\times 8$ to $16\times 16$, as shown in Fig. 1.
Fig. 2: Two example spectrograms sliced from the same crackling & breathing cycle at different time intervals.
@@ 3.1 Experimental Settings
The proposed MVST is evaluated on the largest publicly available respiratory sound dataset, the ICBHI dataset [1], which contains 6,898 respiratory cycles, including 3,642 normal breathing, 1,864 crackling breathing, 886 wheezing breathing and 506 breathing with both crackling & wheezing. The sampling rate is 4kHz, 10kHz or 44.1kHz, with a recording duration between 10 and 90 seconds. Each recording includes an average of seven breathing cycles. Following the official experimental settings in [1, 2], 60% of the patient recordings are randomly selected for training and the rest 40% for testing. Standard evaluation metrics [1] such as specificity (SP), sensitivity (SE) and average score (AS) are used. For the proposed MVST, the cross-entropy loss is used with a learning rate of $5\times 10^{-5}$. The AdamW optimizer is used with a fixed weight decay of $1\times 10^{-6}$. To ensure the convergence and the efficiency of the model, the maximum number of epochs is set to $50$ and the batch size is $8$.
@@ 3.2 Ablation Study
An ablation study is carried out to evaluate the effectiveness of the proposed MVST. For the five compared single-view methods, the spectrograms are divided into patches of the specific size and fed into the ViT for classification. The results are summarized in Table 1. It can be observed that: 1) Many other views perform competitively compared with $16\times 16$ baseline method, which motivates us to jointly learn from different views of different discrimative features. 2) The $128\times 2$ patching scheme produces the highest AS of 62.94% among single-view methods, which demonstrates the effectiveness of splitting into slim and long patches by reducing the time interval while increasing the bandwidth, as this design partially alleviates the problem of frequency misalignment in respiratory sounds. 3) The proposed MVST produces significantly better SP, SE and AS than all single-view methods, which is attributed to the multi-view spectrograms in effectively capturing the audio characteristics, and the gated fusion scheme in combining the multi-view spectrogram features for robust respiratory classification.
@@
$32\times 8$ 79.31 42.61 60.96
$64\times 4$ 84.55 39.68 62.11
$128\times 2$ 80.17 45.71 62.94
$256\times 1$ 84.63 40.14 62.38
@@ 3.3 Comparison to State-of-the-art Methods
Many deep neural networks have been utilized as the backbone for respiratory sound classification, e.g., bi-ResNet in LungBRN [4], ResNet-Att in LungAttn [5], ResNeSt [7], ResNet-34 in RespireNet [6], bi-ResNet-Att in ARSC-Net [3], CNN6 [10], Co-tuned ResNet-50 [2] and audio spectrogram transformer with patch-mix and contrastive learning (AST + patch-mix CL[22]. The comparison with these state-of-the-art methods are summarized in Table 2.111The results of all the compared methods are obtained from here. It can be observed that: 1) Many models utilize traditional CNN models for respiratory classification [2, 4, 5, 6, 3, 7, 10], while the most recent work [22] utilizes audio spectrogram transformer [21] and achieves the previous best results on the ICBHI dataset. This partially motivates us to utilize the ViT as the backbone network. 2) The proposed MVST significantly outperforms all the compared methods in terms of all three criteria SP, SE and AS. Specifically, it outperforms the second-best approach AST + patch-mix CL [22] by 0.33%, 8.09% and 4.18% in terms of SP, SE and AS respectively. Compared to Co-tuned ResNet-50 [2], the proposed MVST achieves a performance gain as high as 8.26% in terms of the key metric AS.